# thebibliography

### v4

@@ Abstract
Effective training of Quantum Neural Networks (QNNs) is vital in the Noisy Intermediate Scale Quantum (NISQ) era, where inherent noise challenges scalability. This paper explores how quantum noise accelerates the onset of barren plateaus (BPs) and examines strategies for mitigating this effect through careful selection of qubit measurement observables. We demonstrate that BPs emerge earlier in noisy environments than in ideal conditions. However, by strategically choosing observables that aligns better with the target objective, QNNs can be trained more effectively. To this end, we study different qubit measurement observables including PauliX, PauliY, PauliZ and a customized Hermitian observable-which aligns better with the learning objective used in this paper. Our analysis covers both global and local cost functions, where global involves all qubits, and local focuses on single-qubit measurements. Our findings reveal that with a global cost function, PauliX and PauliY observables lead to flatter optimization landscapes, indicating BPs as qubit count increases, especially under noise. PauliZ sustains trainability up to $8$ qubits but encounters BPs at $10$ qubits. Notably, the customized Hermitian observable excels in noisy conditions, supporting training up to $10$ qubits. With a local cost function, PauliZ outperforms PauliX and PauliY, maintaining training efficiency up to $10$ qubits. This study underscores the importance of noise consideration in QNN training and offers a strategic approach to observable selection, enhancing QNN performance in noisy quantum environments and advancing quantum machine learning research.
@@ I Introduction
The advent of Noisy Intermediate-Scale Quantum (NISQ) devices—quantum processors with 50 to a few hundred qubits—marks a significant step toward practical quantum computing [1]. By harnessing superposition and entanglement, NISQ devices offer potential advantages in solving classically intractable problems in optimization and simulation [2, 3]. However, their utility is constrained by high noise levels and limited error correction, which become increasingly problematic with scale [1, 4]. Despite these challenges, NISQ devices are already showing promise in cryptography [5], drug discovery [6], financial modeling [7], and quantum machine learning [8, 9].
@@
Several strategies have been proposed to mitigate BPs [15, 16, 17], alongside efforts to understand their underlying causes. Key contributing factors include the nature and degree of entanglement [18], circuit expressibility [19], cost function globality [20, 21], and hardware-induced noise [22]. Among these, quantum noise is particularly critical in the NISQ era, as it is inherent to current quantum devices and directly impacts the trainability of QNNs. Addressing noise-related challenges is thus essential for advancing the practical deployment of QNNs.
@@ I-A Motivational Example
The robustness of QNNs to noise in NISQ devices remains a critical concern. To highlight how noise can induce Barren Plateaus (BPs) and hinder trainability, Fig. 1 illustrates its impact on the cost function landscape. We employ a $4$-qubit QNN composed of alternating $R_{x}(\theta)$ and $R_{y}(\theta)$ rotations, with adjacent qubits entangled via controlled-Z (CZ) gates. The circuit depth is $10$, comprising $80$ single-qubit and 40 two-qubit gates, with measurements in the Pauli-Z basis. The model is trained to learn the identity function using the Adam optimizer (learning rate = $0.1$) for $50$ iterations. Under ideal conditions, the cost landscape (Fig. 1(a)) exhibits a rich structure and multiple minima, enabling efficient optimization. However, in noisy settings where each qubit is affected in every layer, the landscape flattens significantly (Fig. 1(b)), impeding convergence. This degradation is evident in the training performance (Fig. 1(c)), where noise-free training substantially outperforms the noisy counterpart. These findings underscore the importance of addressing noise to realize scalable QNNs. This paper presents a detailed analysis of noise-induced effects, identifying conditions where noise may be leveraged and proposing strategies for its mitigation.
@@ I-B Our Contributions
We investigate how quantum noise and observable choice affect QNN training dynamics and barren plateaus, comparing global vs. local cost functions under fixed-depth circuits with increasing qubits. Our contributions are summarized below:
@@ II-A Qubit Measurement and Observables
This $2\times 2$ Hermitian matrix acts on the two-dimensional complex Hilbert space of a single qubit. In quantum mechanics, $\sigma_{z}$ corresponds to measurements along the Z-axis, distinguishing between the computational basis states |$\ket{0}$ and |$\ket{1}$. Its eigenvalues are $+1$ and $-1$, with $+1$ associated with |$\ket{0}$ and $-1$ with |$\ket{1}$. Thus, a measurement in the PauliZ basis reveals whether the qubit is aligned (|$\ket{0}$) or anti-aligned (|$\ket{1}$) with the Z-axis.
@@
This observable corresponds to measurements along the X-axis in the Bloch sphere representation. When applied to a qubit, it acts as a bit-flip operator, exchanging the computational basis states |$\ket{0}$ and |$\ket{1}$. The PauliX gate is thus analogous to the classical NOT gate, making it fundamental for qubit state manipulation. The eigenvalues of $\sigma_{x}$ are $+1$ and $-1$, but its eigenstates differ from the standard basis. Instead, they are the symmetric and anti-symmetric superposition states:
@@
In quantum mechanics, the PauliY observable corresponds to the measurement of spin or polarization along the Y-axis. When applied to a quantum state, it provides information about the phase relationship between the basis states |$\ket{0}$ and |$\ket{1}$. The eigenvalues of this matrix are $+1$ and $-1$, which are related to these phase-adjusted states. It is similar to the X-gate but also introduces a phase shift. The Y-gate, flips the state of a qubit from |$\ket{0}$ to $i\ket{1}$ and |$\ket{1}$ to $-i\ket{0}$. The eigenstates of the PauliY observable are different from the computational basis states (|$\ket{0}$ and |$\ket{1}$) and are instead specific superpositions of these states with complex coefficients. These states are stable under the action of the PauliY observable.
@@
This implies that for any two vectors |$\ket{a}$ and |$\ket{b}$ in the Hilbert space, the inner product satisfies:
@@ II-B Amplitude Damping
    This operator captures the decay of the qubit from |$\ket{1}$ to |$\ket{0}$ with probability $\gamma$ due to environmental interaction.
@@ III-A Specifications
This corresponds to the projector |$\ket{0000}\bra{0000}$ |, which projects onto the state where all qubits are in the |$\ket{0}$ state. This completely aligns with the training goal (cost function) considered in this paper, which is to learn an Identity gate (1 minus the probability of all qubits being in |$\ket{0}$ state), details of which are presented later in Section IV.
@@ III-B QNN Architecture
    |$\ket{\psi_{final}}=U^{5}\ket{\psi_{0}}$
    where the superscript 5 denotes the number of times the layer is repeated. Finally, the qubits are measured to get the output. For an customized Hermitian operator $M$, the probability of measuring a state |$\ket{\phi}$ is given by:
    $P(\phi)=|\bra{\phi}M\ket{\psi_{final}}|^{2}$
@@
The architecture of local QNN is same as that of global QNN, which is discussed above, however, it differs in terms of qubit measurment, i.e., only a single qubit is measured. For instance, if the $i^{th}$ qubit is measured in PauliZ basis and the outcome is 00, the post-measurement state of the system can be represented as:
|$\ket{\psi^{\prime}}=\frac{(I\otimes\ldots\otimes\bra{0}_{i}\otimes\ldots\otimes I)\ket{\psi_{final}}}{\sqrt{P(0)}}$
where, $P(0)$ is the probability of measuring the qubit in state |$\ket{0}$, and |$\ket{\psi_{final}}$ is the state of the system before measurement and the identity operators $I$ are applied to all qubits except the $i^{th}$ qubit.
@@ IV Experimental Setup
We typically want to maximize the probability of all the qubits being in |$\ket{0}$ state for global cost function and the probability of first qubit in state |$\ket{0}$ for local cost function. This objective closely aligns with the customized Hermitian observable that we have used (See Section III-A) that projects the quantum system to an all |$\ket{0}$ state in case of global cost function. For the training analysis, the QNNs are trained for $50$ training iterations. Adam optimizer with a learning rate of $0.1$ is used for the optimization. The experiments are performed using Pennylane [23].
V Results and Discussion
@@
V-A Global QNN - Comparative Analysis of QNNs Training Dynamics: Noisy vs. Ideal Conditions
@@
V-B Global QNN - Focused Analysis QNNs Training Dynamics in Noisy Setting
@@
V-C Local QNN: Analysis of QNNs Training Dynamics in Noisy Conditions
In this section, we shift our focus to local QNNs, where only a single qubit is measured. Our previous investigations, detailed in Section V-A, demonstrated that a customized Hermitian observable, well-aligned with our chosen cost function (refer to Section IV), outperformed standard observables (PauliZ, PauliX, and PauliY) in global QNN setup. Thus, this analysis will solely consider these standard observables to evaluate whether localizing the cost function enhances the overall performance of QNNs with these observables. Furthermore, our previous results indicated that in the absence of quantum noise, all observables showed high effectiveness, especially in systems with lower qubit counts (4 and 6 qubits). However, under quantum noise and at higher qubit counts (8 and 10 qubits), these observables encountered barren plateaus (BPs) rapidly. Therefore, this analysis will concentrate on noisy conditions at these larger qubit counts.
V-C1 8-Qubit Local QNN
@@ V-C Local QNN: Analysis of QNNs Training Dynamics in Noisy Conditions
Figure 10: Training Results of 8 and 10 Qubit Local QNN with different observables in Noisy Environment. The landscape with PauliZ observable in local QNN even with 10 qubits still have multiple wider regions leading to the solution. PauliX and PauliY observables’ landscape are still completely flat.
@@
This comparative analysis highlights the superior adaptability and learning efficiency of local QNN when the PauliZ observable is implemented, as opposed to the limited training efficacy observed with the PauliX and PauliY observables. Such outcomes underline the critical importance of selecting appropriate measurement observables in QNN configurations to achieve optimal training and performance results.
V-C2 10-Qubit Local QNN
The optimization landscapes in case of 10-qubit local QNN with local cost function design, as shown in Fig. 11, do not differ much to that of 8-qubit QNN (with local cost function) design (Fig. 9). The PauliZ observable still exhibits a pretty dynamic landscape with multiple regions containing the solution making it suitable for the optimizers to find the solution, as shown in label 1 of Fig. 11. This shows a greater potential of PauliZ observable when used local cost function definition, to overcome the so called BPs problem even at higher qubit count and under noisy conditions. On the other hand, the optimization landscape for PauliX and PauliY observables are completely flat and are not suitable for the optimization, as shown in label 2 and 3 in Fig. 11 analogous to the case of 8 qubit QNN with local cost function ( 2 and 3 in Fig. 9).
@@
VI Conclusion
This paper explores the impact of quantum noise on QNN trainability, highlighting how observable selection and cost function design influence the onset of barren plateaus (BPs) in the NISQ era. We show that PauliZ with a global cost function maintains good trainability up to 68686–86 – 8 qubits but fails at $10$, while PauliX and PauliY lead to earlier BPs. Our key finding is that a custom Hermitian observable with global cost not only resists noise-induced degradation but leverages noise to enhance training up to $10$ qubits. PauliZ also performs best under local cost settings. Future work will extend to other noise types and broader problem domains.
@@ Acknowledgements
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008.

### v5

@@ Abstract
Effective training of Quantum Neural Networks (QNNs) is crucial in the Noisy Intermediate-Scale Quantum (NISQ) era, where noise accelerates the onset of barren plateaus (BPs) and limits scalability. This paper investigates how quantum noise impacts QNN trainability and demonstrates that careful selection of qubit measurement observables can mitigate these effects. We analyze PauliX, PauliY, PauliZ, and a customized Hermitian observable under both global (all-qubit measured) and local (single-qubit measured) cost functions. Our results show that with global cost function, PauliX and PauliY lead to flatter landscapes under noise, while PauliZ maintains training up to $8$ qubits before encountering BPs. The customized Hermitian observable proves most robust, enabling training up to $10$ qubits in noisy settings. For local cost function setting, PauliZ outperforms PauliX and PauliY, maintaining efficiency up to $10$ qubits. These findings highlight the importance of noise-aware observable selection, offering a practical strategy to improve QNN performance and advance quantum machine learning in noisy environments.
@@ I Introduction
The advent of Noisy Intermediate-Scale Quantum (NISQ) devices marks a significant step toward practical quantum computing [1]. By harnessing superposition and entanglement, NISQ devices offer potential advantages in solving classically intractable problems in optimization and simulation [2, 3, 4]. However, their utility is constrained by high noise levels and limited error correction, which become increasingly problematic with scale [1, 5, 6]. Despite these challenges, NISQ devices are already showing promise in cryptography [7], drug discovery [8], and quantum machine learning [9, 10, 11].
@@
Several strategies have been proposed to mitigate BPs [17, 18, 19, 20, 21], alongside efforts to understand their underlying causes. Key contributing factors include the nature and degree of entanglement [22], circuit expressibility [23], cost function globality [24, 25], and hardware-induced noise [26, 27]. Among these, quantum noise is particularly critical in the NISQ era, as it is inherent in current quantum devices and directly impacts the trainability of QNNs [28, 29, 26]. Addressing noise-related challenges is thus essential for advancing the practical deployment of QNNs.
@@ I-A Motivational Example
The robustness of QNNs to noise in NISQ devices remains a critical concern. To highlight how noise can give rise to early occurence of BPs leading to reduced trainability, Fig. 1 illustrates its impact on the cost function landscape. We employ a $4$-qubit QNN composed of alternating $R_{x}(\theta)$ and $R_{y}(\theta)$ rotations, with adjacent qubits entangled via controlled-Z (CZ) gates. The circuit depth is $10$, comprising $80$ single-qubit and 40 two-qubit gates, with measurements in the Pauli-Z basis. The model is trained to learn the identity function using the Adam optimizer (learning rate = $0.1$) for $50$ iterations. Under ideal conditions, the cost landscape (Fig. 1(a)) exhibits a rich structure and multiple minima, enabling efficient optimization. However, in noisy settings where each qubit is affected in every layer, the landscape flattens significantly (Fig. 1(b)), impeding convergence. This degradation is evident in the training performance (Fig. 1(c)), where noise-free training substantially outperforms the noisy counterpart. These findings underscore the importance of addressing noise to realize scalable QNNs. This paper presents a detailed analysis of noise-induced effects on QNNs training, identifying conditions where noise may be leveraged and proposing strategies for its mitigation.
@@ I-B Our Contributions
We investigate how quantum noise and observable choice affect QNN training dynamics in global and local cost functions under fixed-depth circuits with increasing qubits. Our contributions are summarized below:
@@ II-A Qubit Measurement and Observables
[…]
@@
[…]
@@
[…]
@@
[…]
@@ II-B Amplitude Damping
[…]
@@ III-A Specifications
[…]
@@ III-B QNN Architecture
[…]
@@
The architecture of local QNN is same as that of global QNN, which is discussed above, however, it differs in terms of qubit measurment, i.e., only a single qubit is measured. For instance, if the $i^{th}$ qubit is measured in PauliZ basis and the outcome is 0, the post-measurement state of the system can be represented as:
[…]
@@ III-E Experimental Setup
[…]
IV Results and Discussion
@@
IV-A Global QNN - Comparative Analysis of QNNs Training Dynamics: Noisy vs. Ideal Conditions
@@
IV-B Global QNN - Focused Analysis QNNs Training Dynamics in Noisy Setting
@@
IV-C Local QNN: Analysis of QNNs Training Dynamics in Noisy Conditions
In this section, we shift our focus to local QNNs, where only a single qubit is measured. Our previous investigations, detailed in Section IV-A, demonstrated that a customized Hermitian observable, well-aligned with our chosen cost function (refer to Section III-E), outperformed standard observables (PauliZ, PauliX, and PauliY) in global QNN setup. Thus, this analysis will solely consider these standard observables to evaluate whether localizing the cost function enhances the overall performance of QNNs with these observables. Furthermore, our previous results indicated that in the absence of quantum noise, all observables showed high effectiveness, especially in systems with lower qubit counts (4 and 6 qubits). However, under quantum noise and at higher qubit counts (8 and 10 qubits), these observables encountered barren plateaus (BPs) rapidly. Therefore, this analysis will concentrate on noisy conditions at these larger qubit counts.
IV-C1 8-Qubit Local QNN
@@ IV-C Local QNN: Analysis of QNNs Training Dynamics in Noisy Conditions
Figure 10: Training Results of 88 and $10$ Qubit Local QNN with different observables in Noisy Environment. The landscape with PauliZ observable in local QNN even with $10$ qubits contains multiple wider regions leading to the solution. PauliX and PauliY observables’ landscape are still completely flat.
@@
This comparative analysis highlights the superior adaptability and learning efficiency of local QNN when the PauliZ observable is implemented, as opposed to the limited training efficacy observed with the PauliX and PauliY observables. Such outcomes underline the critical importance of selecting appropriate measurement observables in QNN configurations to achieve optimal training performance.
IV-C2 10-Qubit Local QNN
The optimization landscapes in case of 10-qubit local QNN with local cost function design, as shown in Fig. 11, do not differ much to that of 8-qubit QNN (with local cost function) design (Fig. 9). The PauliZ observable still exhibits a pretty dynamic landscape with multiple regions containing the solution making it suitable for the optimizers to find the solution, as shown in label 1 of Fig. 11. This shows a greater potential of PauliZ observable when used local cost function definition, to overcome the so called BPs problem even at higher qubit count and under noisy conditions. On the other hand, the optimization landscape for PauliX and PauliY observables are completely flat and are not suitable for the optimization (label 2 and 3 in Fig. 11) analogous to the case of 8 qubit QNN with local cost function ( 2 and 3 in Fig. 9).
@@
V Conclusion
This paper explores the impact of quantum noise on QNN trainability, highlighting how observable selection and cost function design influence the onset of barren plateaus (BPs) in the NISQ era. We show that PauliZ with a global cost function maintains good trainability up to 686–8 qubits but fails at $10$, while PauliX and PauliY lead to earlier BPs. Our key finding is that a custom Hermitian observable with global cost not only resists noise-induced degradation but leverages noise to enhance training up to $10$ qubits. PauliZ also performs best under local cost settings. Future work will extend to other noise types and broader problem domains.
@@ Acknowledgements
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and the Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104.